OpenStudy (anonymous):

PLS HELP WILL GIVE MEDAL AND FAN!!!!!!!!!!!!1 Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 A) He should have substituted 5 + 2x B) He did not distribute correctly C) He combined like terms incorrectly, it should have been 4x instead of 5x D) He made no mistake

OpenStudy (fortytherapper):

Any ideas?

OpenStudy (anonymous):

No......

OpenStudy (fortytherapper):

Check out line 2 and 3 What's -2 times 5

OpenStudy (anonymous):

-10

OpenStudy (fortytherapper):

On the next step, did he put -10

OpenStudy (anonymous):

No he didn't

Directrix (directrix):

x − 2(5 − 2x) = 10 x + 10 + 4x = 10 Does − 2(5 − 2x) = 10 + 4x ?

OpenStudy (faiqraees):

-2 x 5 = -10 not 10

OpenStudy (faiqraees):

So it should be B

OpenStudy (anonymous):

Thanks!!!

OpenStudy (scarlettfarra2000):

If you got the answer please close the question thank you